# Does LLM Focus on the Right Words? Diagnosing Language Bias in LLM-based Recommenders

### v1

@@ Abstract.
Large language models (LLMs), owing to their extensive open-domain knowledge and semantic reasoning capabilities, have been increasingly integrated into recommender systems (RS). However, a substantial gap remains between the pre-training objectives of LLMs and the specific requirements of recommendation tasks. To address this gap, supervised fine-tuning (SFT) is commonly performed on specially curated recommendation datasets to further enhance their predictive ability. Despite its success, SFT exhibits a critical limitation: it induces Language Bias, whereby the model over-relies on auxiliary tokens—such as task descriptions and prefix-generated tokens—while underutilizing core user interaction tokens that encode user-specific preferences. This bias not only undermines recommendation accuracy but also raises unfairness concerns.
To address this issue, we propose Group Distributionally Robust Optimization-based Tuning (GDRT), a novel fine-tuning paradigm that enforces consistent model performance across token groups with varying degrees of relevance to auxiliary tokens. By adaptively upweighting underperforming groups, typically those weakly correlated with auxiliary tokens, GDRT shifts the model’s attention from superficial auxiliary cues to informative user interaction tokens, thereby mitigating language bias. Extensive experiments conducted on three public datasets demonstrate that GDRT effectively mitigates language bias, yielding substantial improvements in recommendation accuracy (with an average NDCG@10 gain of 24.29%) and significantly enhancing recommendation fairness.
@@ 1. Introduction
With remarkable open-domain knowledge and semantic reasoning capabilities (Guo et al., 2025; Achiam et al., 2023), Large Language Models (LLMs) have been extensively explored for integration into recommendation systems (RS) (Wu et al., 2024). One prominent approach involves positioning LLMs as the central recommendation backbone (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025, 2024; Liao et al., 2024b; Kim et al., 2024; Hou et al., 2024; Wang and Lim, 2023; Tan et al., 2024; Zheng et al., 2024a). These methods express items as textual descriptions (e.g., titles), and construct language prompts based on users’ past interactions, which are then used to instruct LLMs to predict users’ future interactions. Figure 1 illustrates the mechanism of such LLM-based recommendation. LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items by analyzing their nuanced semantic relations with the user’s previous interactions and other auxiliary information (e.g., task descriptions, prefix tokens of the prediction). This fine-grained paradigm enables the capture of subtle semantic patterns in user interests and thus represents a promising direction for advancing recommender systems (Bao et al., 2025).
Figure 1. Illustration of LLM‑based recommendations and language bias, wherein the model exhibits an over‑reliance on auxiliary tokens and insufficient utilization of interaction tokens during generation.:
Figure 2. Ratio of attribution values between auxiliary tokens and user interaction tokens before and after SFT. Left: task description vs. user interaction tokens. Right: prefix tokens of predicted item (take the first token) vs. user interaction tokens.:
@@
However, we find that SFT introduces a significant Language Bias. Specifically, SFT drives the model to over-rely on auxiliary tokens (e.g., task descriptions or prefix tokens) while under-utilizing core interaction tokens that encode user personalized preferences as shown in Figure 1. To verify this, we conduct Feature Ablation Attribution analysis (Kokhlikyan et al., 2020; Miglani et al., 2023), a standard approach to quantify each token’s contribution to model predictions. Figure 2 shows that SFT dramatically amplifies the relative impact of auxiliary tokens while significantly suppressing the influence of interaction tokens, with the influence ratio shifting from about 1:1 before fine-tuning to more than 6:1 afterward as measured on typical Amazon datasets.
Figure 3. Distribution of Top‑1 recommended items generated by SFT-trained LLMs across five group defined according to the semantic relevance of items to the auxiliary tokens (Group 1: highest relevance, Group 5: lowest relevance). We also present the distribution of the test set across the same five groups for comparison.:
This over-reliance reveals shortcut learning: LLMs simply memorize correlations with frequently occurring auxiliary tokens rather than grounding predictions in user-specific preferences. Such bias not only undermines recommendation accuracy but also raises serious fairness concerns (Chen et al., 2023). Specifically, this bias skews recommendations toward a narrow subset of items whose tokens exhibit higher semantic relevance to auxiliary tokens. To verify this, we divided items into five groups based on their relevance to auxiliary tokens. As shown in Figure 3, over 80% of recommended items fall into the most relevant group, while this group comprises only 20% of target items in the test set. These findings motivate our core research question: How can we mitigate language bias in LLM-based recommenders?
To address this, we propose a novel fine-tuning strategy GDRT, that leverages Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) to mitigate language bias. We first group training instances according to the semantic relevance between the target token and the auxiliary tokens, which can be evaluated by the LLM’s predictive probability when user history is masked. Group DRO then enforces LLMs to perform consistently well across all groups, regardless of their relevance strength with auxiliary tokens. This optimization objective naturally shifts the LLM’s attention away from auxiliary tokens toward user-specific interaction tokens, as simple reliance on shortcut auxiliary tokens results in poor performance on groups with weaker correlations. Importantly, GDRT is easy to implement and computationally efficient, requiring only high‑efficiency group construction and dynamic sample weighting during training. It can be seamlessly integrated into various LLM-based recommenders, yielding improvements in both accuracy and fairness.
@@
  • We provide a comprehensive empirical analysis revealing that SFT in LLM-based recommendation induces significant language bias, negatively affecting both accuracy and fairness.
  • We propose GDRT, a Group DRO-based fine-tuning strategy, to effectively mitigate language bias in LLM-based recommendation.
@@ 2.1. LLM-based Recommendation
The remarkable success of LLMs across diverse domains (Wu et al., 2023b; Ouyang et al., 2022; Tang et al., 2024) has spurred growing interest in their application to recommendation systems (RS) (Wu et al., 2024). A prominent approach is to directly leverage LLMs as recommenders (Li et al., 2023c). As shown in Figure 1, this paradigm constructs a language prompt $x=[\,x^{\text{task}};\,x^{\text{user}}\,]$, where $x^{\text{task}}$ represents the task description and $x^{\text{user}}$ denotes the textual form (e.g., titles) of a user’s historical interactions. This prompt then guides the LLM to generate the descriptions of recommended items $y$. Notably, LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items according to the model estimated probability $P_{\theta}(y_{t}|x,y_{<t})$, where $y_{t}$ denotes the $t$-th predictive tokens and $y_{<t}$ denotes the prefix tokens of the prediction. This fine-grained token-level paradigm has the potential to capture subtle semantic patterns in user preferences.
@@
2.2. Analyses on Language Bias
In this section, we first identify the language bias in fine-tuning LLM for recommendation, followed by discussing its negative effect. We then analyze the underlying causes of this bias and discuss why existing methods can not effectively address this issue.
2.2.1. Empirical Evidence Demonstrating Language Bias
@@
Language Bias: Supervised fine-tuning (SFT) can bias LLMs to over-rely on auxiliary tokens while under-utilizing core interaction tokens that encode user personalized preferences.
Before fine-tuning, the attribution value ratio between task description and interaction tokens is approximately 1:1. After fine-tuning, this ratio is markedly amplified across all datasets (e.g., Toy: 4.69:1; Clothing: 6.58:1). A similar phenomenon is observed when comparing prefix tokens with interaction tokens. These results clearly indicate the presence of language bias, whereby the model disproportionately relies on auxiliary tokens rather than more informative interaction tokens.
Figure 4. Proportion of Top‑1 recommendations belonging to Item Group 1 (most relevant with auxiliary tokens) over the course of SFT.:
2.2.2. Negative Effect of Language Bias
Language bias can substantially hinder the effectiveness of LLM-based recommenders, leading not only to decreased recommendation accuracy but also to pronounced unfairness issues. On the one hand, critical user–item interaction signals that capture user preferences may be ignored by the model, severely impairing its ability to deliver personalized recommendations. On the other hand, the recommendation output becomes inherently skewed toward a limited subset of items whose textual tokens exhibit strong correlations with auxiliary tokens.
@@ 2.2.2. Negative Effect of Language Bias
Here we mask the user historical information in the original prompt template with the placeholder ’N/A’, and compute the probability of the tokens for each item. This measure serves as an indicator of the relevance between the input and the output, a practice commonly adopted for estimating semantic relevance (Kauf et al., 2024). Based on this metric, items are sorted and evenly divided into five groups, where Group 1 contains the most relevant items, and Group 5 the least.
@@
Figure 5. The co-occurrence rate of different types of token pairs in the training set.:
2.2.3. Origins of Language Bias
The emergence of language bias can be traced to biases inherent in the training data. As shown in Figure 5, the co-occurrence rate between auxiliary tokens and target item tokens is significantly higher than those between interaction tokens and target item tokens. This is expected, as task prompts appear in every training instance, and prefix tokens are always accompanied for item tokens. Consequently, during fine-tuning, the LLM tends to capture shortcut patterns, simply memorizing frequent correlations with auxiliary tokens, while neglecting the more important user-specific interaction signals. This incurs language bias and motivates the need for improved fine-tuning strategies.
@@ 2.2.4. Limitations of Related Strategies
Several recent studies have explored bias and fairness issues in LLM-based RS. However, the characteristics of these biases differ from those of language bias. Specifically, language bias arises during fine‑tuning and reflects the inherent over‑reliance of LLMs on auxiliary tokens. In contrast, earlier work has primarily focused on biases such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a), arising from imbalanced item frequency in training data; position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), caused by the model’s sensitivity to the ordering of candidate items; length bias (Bao et al., 2024), induced by length normalization, which makes items with longer token sequences more likely to be recommended. These biases stem from distinct sources and mechanisms. Consequently, this newly identified form of language bias warrants explicit and targeted mitigation, yet recent work has fallen short in addressing this issue (cf.  Table 2).
Another related work, CFT (Zhang et al., 2024b), aims to encourage LLMs to better leverage user interaction information. Notably, this study is parallel to ours, currently available only on arXiv and not formally published. CFT introduces counterfactual learning to forcibly enhance the influence of interaction tokens for each training instance. However, this strategy suffers from multiple aspects of limitations: (1) Objective misalignment. The counterfactual objective is not directly aligned with improving recommendation accuracy or fairness, and its contribution to these aspects remains uncertain. In fact, CFT only treats this objective as an auxiliary loss and over-emphasis on this term has been observed to lead to substantial performance degradation. (2) Difficulty in weight selection. CFT relies on manually specified weights to determine the degree to which a training instance should rely on interaction tokens. In fact, the optimal weights can vary substantially across instances and are difficult to estimate accurately. Manual specification is therefore challenging, prone to deviation from the ideal value, and ultimately detrimental to model effectiveness. Empirically, even when we directly employ the official source code of CFT and conduct a fine‑grained hyperparameter search, CFT yields only limited performance gains on some datasets (cf.  Table 2). (3) High computational overhead. CFT requires to process counterfactual instances, resulting in more than double the training time compared with SFT (cf. Figure 12). In contrast, our GDRT is explicitly aligned with the target objective, enabling the model to consistently perform well across different instance groups. This naturally encourages greater focus on interaction tokens without requiring manual specification of influence strengths.
@@ 3. Methodology
The above analyses reveal a significant language bias inherent in fine-tuning LLMs for recommendation. To address this issue, we propose a novel fine‑tuning framework, termed GDRT, which leverages Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) to mitigate such bias. This section first outlines the general idea of GDRT, and then describes the group partitioning strategy and customized loss function.
General Idea. Rather than directly intervening the learning process to manually enhance the effect of interaction tokens, we pursue an alternative objective: ensuring that the model achieves consistently strong performance across target tokens regardless of their degree of relevance to auxiliary tokens. It naturally shifts the LLM’s attention away from auxiliary tokens towards user‑specific interaction tokens, as simple reliance on shortcut auxiliary tokens results in poor performance on the instances with weak correlations as shown in Figure 6. Besides, this objective is directly aligned with the goals of high recommendation accuracy and fairness.
@@
Figure 6. Loss of token groups with varying degrees of relevance to auxiliary tokens in the SFT-trained biased model.:

### v2

# Does LLM Focus on the Right Words? Mitigating Context Bias in LLM-based Recommenders

@@ Abstract.
Large language models (LLMs), owing to their extensive open-domain knowledge and semantic reasoning capabilities, have been increasingly integrated into recommender systems (RS). However, a substantial gap remains between the pre-training objectives of LLMs and the specific requirements of recommendation tasks. To address this gap, supervised fine-tuning (SFT) is commonly performed on specially curated recommendation datasets to further enhance their predictive ability. Despite its success, SFT exhibits a critical limitation: it induces Context Bias, whereby the model over-relies on auxiliary tokens—such as task descriptions and prefix-generated tokens—while underutilizing core user interaction tokens that encode user-specific preferences. This bias not only undermines recommendation accuracy but also raises unfairness concerns.
To address this issue, we propose Group Distributionally Robust Optimization-based Tuning (GDRT), a novel fine-tuning paradigm that enforces consistent model performance across token groups with varying degrees of relevance to auxiliary tokens. By adaptively upweighting underperforming groups, typically those weakly correlated with auxiliary tokens, GDRT shifts the model’s attention from superficial auxiliary cues to informative user interaction tokens, thereby mitigating context bias. Extensive experiments conducted on three public datasets demonstrate that GDRT effectively mitigates context bias, yielding substantial improvements in recommendation accuracy (with an average NDCG@10 gain of 24.29%) and significantly enhancing recommendation fairness. The code is available at https://github.com/WANGBohaO-jpg/GDRT.
@@ 1. Introduction
With remarkable open-domain knowledge and semantic reasoning capabilities (Guo et al., 2025; Achiam et al., 2023), Large Language Models (LLMs) have been extensively explored for integration into recommendation systems (RS) (Wu et al., 2024). One prominent approach involves positioning LLMs as the central recommendation backbone (Li et al., 2023c; Bao et al., 2025, 2023; Zhu et al., 2024; Wang et al., 2025b, 2024b; Liao et al., 2024b; Kim et al., 2024; Hou et al., 2024; Wang and Lim, 2023; Tan et al., 2024; Zheng et al., 2024a). These methods express items as textual descriptions (e.g., titles), and construct language prompts based on users’ past interactions, which are then used to instruct LLMs to predict users’ future interactions. Figure 1 illustrates the mechanism of such LLM-based recommendation. LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items by analyzing their nuanced semantic relations with the user’s previous interactions and other auxiliary information (e.g., task descriptions, prefix tokens of the prediction). This fine-grained paradigm enables the capture of subtle semantic patterns in user interests and thus represents a promising direction for advancing recommender systems (Bao et al., 2025).
Figure 1. Illustration of LLM‑based recommendations and context bias, wherein the model exhibits an over‑reliance on auxiliary tokens and insufficient utilization of interaction tokens during generation.
Figure 2. Ratio of attribution values between auxiliary tokens and user interaction tokens before and after SFT. Left: task description vs. user interaction tokens. Right: prefix tokens of predicted item (take the first token) vs. user interaction tokens.
@@
However, we find that SFT introduces a significant Context Bias. Specifically, SFT drives the model to over-rely on auxiliary tokens (e.g., task descriptions or prefix tokens) while under-utilizing core interaction tokens that encode user personalized preferences as shown in Figure 1. To verify this, we conduct Feature Ablation Attribution analysis (Kokhlikyan et al., 2020; Miglani et al., 2023), a standard approach to quantify each token’s contribution to model predictions. Figure 2 shows that SFT dramatically amplifies the relative impact of auxiliary tokens while significantly suppressing the influence of interaction tokens, with the influence ratio shifting from about 1:1 before fine-tuning to more than 6:1 afterward as measured on typical Amazon datasets.
Figure 3. Distribution of Top‑1 recommended items generated by SFT-trained LLMs across five group defined according to the semantic relevance of items to the auxiliary tokens (Group 1: highest relevance, Group 5: lowest relevance). We also present the distribution of the test set across the same five groups for comparison.
This over-reliance reveals shortcut learning: LLMs simply memorize correlations with frequently occurring auxiliary tokens rather than grounding predictions in user-specific preferences. Such bias not only undermines recommendation accuracy but also raises serious fairness concerns. Specifically, this bias skews recommendations toward a narrow subset of items whose tokens exhibit higher semantic relevance to auxiliary tokens. To verify this, we divided items into five groups based on their relevance to auxiliary tokens. As shown in Figure 3, over 80% of recommended items fall into the highest relevance group, while this group comprises only 20% of target items in the test set. These findings motivate our core research question: How can we mitigate context bias in LLM-based recommenders?
To address this, we propose a novel fine-tuning strategy GDRT, that leverages Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) to mitigate context bias. We first group training instances according to the semantic relevance between the target token and the auxiliary tokens, which can be evaluated by the LLM’s predictive probability when user history is masked. Group DRO then enforces LLMs to perform consistently well across all groups, regardless of their relevance strength with auxiliary tokens. This optimization objective naturally shifts the LLM’s attention away from auxiliary tokens toward user-specific interaction tokens, as simple reliance on shortcut auxiliary tokens results in poor performance on groups with weaker correlations. Importantly, GDRT is easy to implement and computationally efficient, requiring only high‑efficiency group construction and dynamic sample weighting during training. It can be seamlessly integrated into various LLM-based recommenders, yielding improvements in both accuracy and fairness.
@@
  • We provide a comprehensive empirical analysis revealing that SFT in LLM-based recommendation induces significant context bias, negatively affecting both accuracy and fairness.
  • We propose GDRT, a Group DRO-based fine-tuning strategy, to effectively mitigate context bias in LLM-based recommendation.
@@ 2.1. LLM-based Recommendation
The remarkable success of LLMs across diverse domains (Wu et al., 2023b; Ouyang et al., 2022; Tang et al., 2024; Chen et al., 2025c; Wang et al., 2025a; Chen et al., 2025b) has spurred growing interest in their application to recommendation systems (RS) (Wu et al., 2024). A prominent approach is to directly leverage LLMs as recommenders (Li et al., 2023c). As shown in Figure 1, this paradigm constructs a language prompt $x=[\,x^{\text{task}};\,x^{\text{user}}\,]$, where $x^{\text{task}}$ represents the task description and $x^{\text{user}}$ denotes the textual form (e.g., titles) of a user’s historical interactions. This prompt then guides the LLM to generate the descriptions of recommended items $y$. Notably, LLMs operate at a fine-grained semantic token level, sequentially generating tokens of the predicted items according to the model estimated probability $P_{\theta}(y_{t}|x,y_{<t})$, where $y_{t}$ denotes the $t$-th predictive tokens and $y_{<t}$ denotes the prefix tokens of the prediction. This fine-grained token-level paradigm has the potential to capture subtle semantic patterns in user preferences.
@@
2.2. Analyses on Context Bias
In this section, we first identify the context bias in fine-tuning LLM for recommendation, followed by discussing its negative effect. We then analyze the underlying causes of this bias and discuss why existing methods can not effectively address this issue.  [^1] The configuration of the empirical analysis in this section is consistent with our main experimental setup in Section 4.1.3. Additional experiments on other Prompts and LLMs are provided in the Appendix A.1.
2.2.1. Empirical Evidence Demonstrating Context Bias
@@
Context Bias: Supervised fine-tuning (SFT) can bias LLMs to over-rely on auxiliary tokens while under-utilizing core interaction tokens that encode user personalized preferences.
Before fine-tuning, the attribution value ratio between task description and interaction tokens is approximately 1:1. After fine-tuning, this ratio is markedly amplified across all datasets (e.g., Toy: 4.69:1; Clothing: 6.58:1). A similar phenomenon is observed when comparing prefix tokens with interaction tokens. These results clearly indicate the presence of context bias, whereby the model disproportionately relies on auxiliary tokens rather than more informative interaction tokens.
Figure 4. Proportion of Top‑1 recommendations belonging to Item Group 1 (highest relevance with auxiliary tokens) over the course of SFT.
2.2.2. Negative Effect of Context Bias
Context bias can substantially hinder the effectiveness of LLM-based recommenders, leading not only to decreased recommendation accuracy but also to pronounced unfairness issues. On the one hand, critical user–item interaction signals that capture user preferences may be ignored by the model, severely impairing its ability to deliver personalized recommendations. On the other hand, the recommendation output becomes inherently skewed toward a limited subset of items whose textual tokens exhibit strong correlations with auxiliary tokens.
@@ 2.2.2. Negative Effect of Context Bias
Here we mask the user historical information in the original prompt template with the placeholder ’N/A’, and compute the probability of the tokens for each item. This measure serves as an indicator of the relevance between the input and output, a practice commonly adopted for estimating semantic relevance (Kauf et al., 2024). Based on this metric, items are sorted and evenly divided into five groups, with Group 1 containing the items of highest relevance and Group 5 the least.
@@
Figure 5. The co-occurrence rate of different types of token pairs in the training set.
2.2.3. Origins of Context Bias
The emergence of context bias can be traced to biases inherent in the training data. As shown in Figure 5, the co-occurrence rate between auxiliary tokens and target item tokens is significantly higher than those between interaction tokens and target item tokens. This is expected, as task prompts appear in every training instance, and prefix tokens are always accompanied for item tokens. Consequently, during fine-tuning, the LLM tends to capture shortcut patterns, simply memorizing frequent correlations with auxiliary tokens, while neglecting the more important user-specific interaction signals. This incurs context bias and motivates the need for improved fine-tuning strategies.
@@ 2.2.4. Limitations of Related Strategies
Several recent studies have explored bias and fairness issues in LLM-based RS. However, the characteristics of these biases differ from those of context bias. Specifically, context bias arises during fine‑tuning and reflects the inherent over‑reliance of LLMs on auxiliary tokens. In contrast, earlier work has primarily focused on biases such as popularity bias (Gao et al., 2025b; Lu et al., 2025; Jiang et al., 2024; Gao et al., 2025a; Liao et al., 2024a), arising from imbalanced item frequency in training data; position bias (Chen et al., 2023; Ma et al., 2023; Hou et al., 2024; Bito et al., 2025; Jiang et al., 2025; Dai et al., 2024; Luo et al., 2025; Zhang et al., 2024a), caused by the model’s sensitivity to the ordering of candidate items; amplification bias (Bao et al., 2024), arising from length normalization, which favors items containing tokens with generation probabilities close to 1. These biases stem from distinct sources and mechanisms. Consequently, this newly identified form of context bias warrants explicit and targeted mitigation, yet recent work has fallen short in addressing this issue (cf.  Table 2).
Another related work, CFT (Zhang et al., 2024b), aims to encourage LLMs to better leverage user interaction information. CFT introduces counterfactual learning to forcibly enhance the influence of interaction tokens for each training instance. However, this strategy suffers from multiple aspects of limitations: (1) Objective misalignment. The counterfactual objective is not directly aligned with improving recommendation accuracy or fairness, and its contribution to these aspects remains uncertain. In fact, CFT only treats this objective as an auxiliary loss and over-emphasis on this term has been observed to lead to substantial performance degradation. (2) Difficulty in weight selection. CFT relies on manually specified weights to determine the degree to which a training instance should rely on interaction tokens. In fact, the optimal weights can vary substantially across instances and are difficult to estimate accurately. Manual specification is therefore challenging, prone to deviation from the ideal value, and ultimately detrimental to model effectiveness. Empirically, even when we directly employ the official source code of CFT and conduct a fine‑grained hyperparameter search, CFT yields only limited performance gains on some datasets (cf.  Table 2). (3) High computational overhead. CFT requires to process counterfactual instances, resulting in more than double the training time compared with SFT (cf. Figure 12). In contrast, our GDRT is explicitly aligned with the target objective, enabling the model to consistently perform well across different instance groups. This naturally encourages greater focus on interaction tokens without requiring manual specification of influence strengths.
@@ 3. Methodology
The above analyses reveal a significant context bias inherent in fine-tuning LLMs for recommendation. To address this issue, we propose a novel fine‑tuning framework, termed GDRT, which leverages Group Distributionally Robust Optimization (Group DRO) (Sagawa et al., 2019) to mitigate such bias. This section first outlines the general idea of GDRT, and then describes the group partitioning strategy and customized loss function.
General Idea. Rather than directly intervening the learning process to manually enhance the effect of interaction tokens, we pursue an alternative objective: ensuring that the model achieves consistently strong performance across target tokens regardless of their degree of relevance to auxiliary tokens. It naturally shifts the LLM’s attention away from auxiliary tokens towards user‑specific interaction tokens, as simple reliance on shortcut auxiliary tokens results in poor performance on the instances with weak correlations. Besides, this objective is directly aligned with the goals of high recommendation accuracy and fairness.
@@
Figure 6. Loss of token groups with varying degrees of relevance to auxiliary tokens in the SFT-trained biased model. Target tokens are grouped into strong- and weak-relevance sets using K-means on relevance scores computed with Eq. 3.